OpenStudy (anonymous):

Use logarithmic differentiation to find the derivative of the function. y= x^(ln(7x))

OpenStudy (anonymous):

log base x (y) = ln(7x) crap... haha, i can't remember what to do next

zepdrix (zepdrix):

\[\Large y=x^{\ln(7x)}\]Taking the natural log of both sides,\[\Large \ln y=\ln [x^{\ln(7x)}]\] Applying a rule of logs:\[\Large \color{teal}{\log(a^b)\quad=\quad b\cdot \log(a)}\]gives us,\[\Large \ln y=\ln(7x)\ln[x]\]

zepdrix (zepdrix):

Confused by any of that? :o

OpenStudy (anonymous):

I understand up to that but I got stuck when I tried to continue.

zepdrix (zepdrix):

So now we need to take the derivative of each side, with respect to x,\[\Large \color{royalblue}{\left(\ln y\right)'} \quad=\quad \color{royalblue}{\left(\ln7x\right)'}\ln x+\ln7x\color{royalblue}{\left(\ln x\right)'}\] Applying the product rule to the right side. So here is our product rule setup. All of the blue parts represent derivatives we need to take.

OpenStudy (anonymous):

alright, so that is: \[y/y' = \ln x/x + \ln 7x/x\] right?

zepdrix (zepdrix):

mmm yah looks good so far!

zepdrix (zepdrix):

woops, left side should be y'/y

OpenStudy (anonymous):

oh yea! Thank you :)

zepdrix (zepdrix):

got it from there? c:

zepdrix (zepdrix):

Oh and \(\Large \color{royalblue}{\text{Welcome to OpenStudy! :)}}\)